# Do Language Models Have a Critical Period for Language Acquisition?

### v1

@@ 7 Discussion
We first address our four research questions from §4. We then explore our results’ implications for the theoretical claims introduced in §3. We close by discussing the limitations of the present study and opportunities for future work.
@@ 7.1 Research Questions
We find evidence against a CP for $\mathrm{L}_{2}$ learning in language models. In experiments 1 and 3 (Figs. 2 and 4), we consistently find that final performance in $\mathrm{L}_{2}$ is worse (or at least not different) in the interleaved than in the sequential condition. This is the opposite of the pattern found in humans, where bilingual learners have greater proficiency in $\mathrm{L}_{2}$ with earlier exposure to it (Johnson and Newport, 1989). Comparing the interleaved and monolingual conditions also informs this question. For humans, early $\mathrm{L}_{2}$ learners resemble monolinguals in their $\mathrm{L}_{2}$ performance. However, monolingual models outperform interleaved ones on BLiMP, and sometimes on GLUE. Looking at the entire learning trajectory, interleaved models differ markedly from both monolingual and sequential models, showing more gradual and delayed improvements.
@@
This RQ concerned the CP for $\mathrm{L}_{1}$ attrition. Humans show few signs of $\mathrm{L}_{1}$ attrition after the CP, even if $\mathrm{L}_{1}$ exposure decreases or ceases. Our results in experiment 2 provide strong evidence against any such effects in typical LMs. Instead, we find that $\mathrm{L}_{1}$ performance worsens rapidly and to a large degree in the sequential condition after $\mathrm{L}_{1}$ exposure ceases. The loss of $\mathrm{L}_{1}$ proficiency is prevented by continuing $\mathrm{L}_{1}$ exposure in the sequential-interleaved condition, but such continued exposure is not necessary in humans to prevent attrition.
@@
This RQ concerned the trade-off between $\mathrm{L}_{2}$ learning and preventing $\mathrm{L}_{1}$ attrition when explicitly reducing plasticity partway during learning. We find strong evidence for such a trade-off when comparing the sequential-EWC condition to the sequential condition. The value of $\lambda$ we selected preserved $\mathrm{L}_{1}$ performance substantially compared to the sequential models, at the cost of harming final performance in $\mathrm{L}_{2}$. The $\mathrm{L}_{2}$ learning curves also converge relatively quickly when plasticity is reduced. Our exploration of different values of $\lambda$ (Fig. 5) shows that preserving $\mathrm{L}_{1}$ performance harms $\mathrm{L}_{2}$ acquisition. Although there is a point for each model architecture where increasing $\lambda$ no longer results in substantially greater $\mathrm{L}_{1}$ preservation, these values of $\lambda$ are too large to allow for monolingual-level $\mathrm{L}_{2}$ performance.
@@
Our results showed that neither language family nor script has a consistent impact on CP effects, though we only look at $\mathrm{L}_{2}$ performance.
@@ 7.2 Theoretical Implications
Our results show that CP effects are not naturally arising in LMs trained in a typical training regime. At the same time, we are able to suggest a methodology to reverse engineer human-like learning patterns by artificially reducing plasticity later in training. As discussed in Section 3, results like these are relevant to certain specific claims in the critical period literature. Specifically, they refute the strong experiential claim, which states that all successful learning algorithms will show CP effects, and they are consistent with (but do not provide strong evidence for) the strong innate claim, which states that innate maturational stages are necessary to produce CP effects.
The strong experiential claim incorrectly predicts that our LMs will naturally show CP effects. This may come as a surprise, given this claim was based on previous studies on connectionist models finding evidence for the phenomenon of entrenchment (Munro, 1986; Ellis and Lambon Ralph, 2000). One explanations for this discrepancy in results is the difference in models’ capacity. These earlier works trained extremely small models by today’s standards, whereas our LMs may be over-parameterized and therefore have sufficient capacity to train to convergence without entrenchment.
On the other hand, the strong innate claim correctly predicts that our LMs will not show CP effects, unless we introduce an innate loss in plasticity. Our introduction of EWC part-way through training is akin to an innate loss in plasticity. However, our experiments are not strong evidence either that strong innate claim is correct or that EWC is a plausible algorithmic mechanism underlying the loss in plasticity in humans. The possibilities remain that other statistical learners will show CP effects as a natural consequence of experience, and that humans could be one such learner.
Finally, many scholars of human development advocate for a complex explanation of CP effects with multiple mechanisms (e.g.,  Newport, 1990; Thiessen et al., 2016; Singleton, 2005). Innate and experiential mechanisms can coexist in humans. We consider this nuanced view to be almost certainly correct, but our results suggest that the role of statistical learning should not be assumed or overstated without evidence from humans or more cognitively plausible models.
@@ 7.3 Limitations and Future Work
From a Bayesian epistemological point of view, results from GPT-2 and RoBERTa should shift our priors about general learners and humans only slightly and in proportion to how representative we consider these models to be of other language learners. More generalizable evidence can be obtained by considering more cognitively-inspired models and learning algorithms, and learning environments drawing on more developmentally plausible linguistic data and multimodal input (Warstadt and Bowman, 2022). Our experiments make a compromise between these goals and maximizing human-like learning outcomes given today’s resources, but future work should revisit this compromise as better resources for cognitive modeling are developed.
One caveat to the rejection of strong experiential claim is that the claim only applies to learners that adequately acquire $\mathrm{L}_{1}$. While this condition is not completely well-defined, it is open to debate whether our models should qualify. For example, our GPT-2 models in the sequential and interleaved conditions achieve $75\pm 1\%$ accuracy on BLiMP. This is substantially better than chance (50%) and is comparable to the strongest baseline model from the BabyLM Challenge, but falls well below human agreement with BLiMP (89%). No LMs currently achieve fully human-like performance on BLiMP, and higher performance generally requires compromising on developmental plausibility, though more data-efficient architectures exist (Samuel et al., 2023; Warstadt et al., 2023). We do not expect our results to be qualitatively different if our experiments are run with more effective Transformer-based LMs, but this is something future work should confirm.
@@
We identify several additional avenues for future work: First, the regularization method we use, EWC (Kirkpatrick et al., 2017), is a computational-level model which induces a predetermined loss in plasticity; future work could consider other regularizers such as Memory Aware Synapsis (Aljundi et al., 2018) that arguably model what happens in humans at the algorithmic-level. Second, there is neurolinguistic evidence for some degree of modularity in how human bilinguals process different languages (Hernandez et al., 2005), but our models use completely shared parameters for languages. Future work can explore LM architectures that encourage or directly build in modularity, such as XLM (Conneau and Lample, 2019) or X-MOD (Pfeiffer et al., 2022). Third, our models learn in a text-only environment, but for humans $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ are both grounded in the same non-linguistic stimuli. Training multimodal models can lead to more realistic simulations, as well as enable testing of CP effects for $\mathrm{L}_{1}$ learning, which requires non-linguistic experience to precede $\mathrm{L}_{1}$ acquisition.
@@ Appendix A Data Preprocessing
We remove extra spaces, non-breaking spaces and dataset-specific characters such as dialogue lines and music symbols for OpenSubtitles or paragraph delimiters for Gutenberg. For the Java code extracted from The Stack, we remove all docstrings and comments.
@@
Interleaved datasets are created for each language pair ($\mathtt{en}$ plus a second language). The same blocks of texts sampled in the previous unifying step are simply interleaved while maintaining their ordering (e.g. $\mathrm{L}_{1}$ block 1, $\mathrm{L}_{2}$ block 1, $\mathrm{L}_{1}$ block 2, $\mathrm{L}_{2}$ block 2, etc.). In this way, the data from the two languages is presented in the same order to the model during the interleaved training as it is during the sequential training.
@@
We uniformize the training dataset sizes across languages. We quantify dataset sizes in terms of the number of tokens obtained using a BPE tokenizer (see section App. A). Beyond the fact that tokens are the true unit of input to the LMs, BPE is a compression algorithm, so while the amount of information per word might be highly language-specific, the amount of information per token is more comparable across languages.
@@ Appendix C Hyperparameters
Table 4: Configurations for GPT-2 models
@@
Table 5: Configurations for RoBERTa models
@@
Table 6: Fixed hyperparameters for GPT-2
@@ Appendix D Evaluation Tasks
Irregular Forms Aaron broke the unicycle. Aaron broken the unicycle
@@
Ellipsis Jill hides one orange chair
  and Tammy hides more.
Jill hides one chair and
  Tammy hides more orange.
@@
The BLiMP tasks that our models are evaluated on are examplified in Tab. 7. The subset of GLUE  [^12] https://gluebenchmark.com tasks that our models are evaluated on is:
@@
Corpus of Linguistic Acceptability consists of around 10K English sentences gathered from published linguistics literature that have been annotated by experts for binary acceptability (grammaticality) judgements (Warstadt et al., 2018).
@@
Binary Stanford Sentiment Treebank includes 215K unique phrases extracted from the parse trees of movie reviews sentences and that have been fully labeled as having either a positive or negative sentiment by three human judges (Socher et al., 2013).
@@
Microsoft Research Paraphrase Corpus contains 5,800 pairs of sentences sourced from a large corpus of news data, each labeled with a binary judgment indicating whether the pair represents a paraphrase or not (Dolan and Brockett, 2005).
@@
Quora Question Pairs is a corpus of over 400K question pairs from the Quora website which are annotated with a binary value denoting whether the questions are paraphrases of each other.
@@
Multi-Genre Natural Language Inference is a crowd-sourced collection of 433K sentence pairs annotated with textual entailment information, covering a wide range of genres of both spoken and written text (Williams et al., 2018).
@@
(Multi-Genre Natural Language Inference Mismatched) is the mismatched version of the MNLI (matched) dataset, where the dev and test sets use out-of-domain data that does not closely resemble anything seen at training time (Williams et al., 2018).
@@
Question-answering Natural Language Inference is a dataset that was automatically derived from the Stanford Question Answering Dataset v1.1 (SQuAD) (Rajpurkar et al., 2016). It consists of question-paragraph pairs with one sentence in each paragraph, sourced from Wikipedia, containing the answer to the corresponding question, written by an annotator (Wang et al., 2019).
@@
Recognizing Textual Entailment is a dataset that was compiled from a series of textual entailment challenges: RTE1 (Dagan et al., 2006), RTE2 (Bar Haim et al., 2006), RTE3 (Giampiccolo et al., 2007), and RTE5 (Bentivogli et al., 2009). The task requires to recognize whether the meaning of a text fragment can be inferred from the other text.
The subset of SuperGLUE  [^13] https://super.gluebenchmark.com tasks that the models are evaluated on is:
@@
Boolean Questions is a reading comprehension dataset consisting of almost 16K naturally occurring yes/no questions that are generated in unprompted and unconstrained settings (Clark et al., 2019).
@@
Multi-Sentence Reading Comprehension contains around 10K questions that can be answered by combining information from a multi-sentence paragraph (Khashabi et al., 2018).
@@
Winograd Schema Challenge is a corpus of sentence pairs that differ in only one or two words and contain an ambiguity that can be resolved using world knowledge and reasoning (Levesque et al., 2011).
Appendix E Derivation of EWC Regularization
We assume we have access to a collection of strings $\mathcal{D}_{\mathrm{L}_{1}}=\{\boldsymbol{x}_{n}\}_{n=1}^{N}$ in $\mathrm{L}_{1}$ and another $\mathcal{D}_{\mathrm{L}_{2}}=\{\boldsymbol{y}_{m}\}_{m=1}^{M}$ in $\mathrm{L}_{2}$. Further, let $\Theta\subset\mathbb{R}^{d}$ be a compact set of possible parameters. We take a Bayesian approach and construct a posterior density over possible parameter vectors $\boldsymbol{\theta}$. Let ${\pi}(\boldsymbol{\theta})$ be a prior density over $\Theta$, and then consider the following likelihood
$p(\mathcal{D}_{\mathrm{L}_{1}},\mathcal{D}_{\mathrm{L}_{2}}\mid\boldsymbol{\theta})=\prod_{n=1}^{N}p(\boldsymbol{x}_{n}\mid\boldsymbol{\theta})\prod_{m=1}^{M}p(\boldsymbol{y}_{m}\mid\boldsymbol{\theta}),$ (2)

### v2

@@ 7 Discussion
We first address our four research questions from §4. We then explore our results’ implications for the theoretical claims introduced in §3.
@@ 7.1 Research Questions
This RQ concerned $\mathrm{L}_{2}$ CP effects. We find evidence against a CP for $\mathrm{L}_{2}$ learning in language models. In experiments 1 and 3 (Figs. 2 and 4), we consistently find that final performance in $\mathrm{L}_{2}$ is worse (or at least not different) in the interleaved than in the sequential condition. This is the opposite of the pattern found in humans, where bilingual learners have greater proficiency in $\mathrm{L}_{2}$ with earlier exposure to it (Johnson and Newport, 1989). Comparing the interleaved and monolingual conditions also informs this question. For humans, early $\mathrm{L}_{2}$ learners resemble monolinguals in their $\mathrm{L}_{2}$ performance. However, monolingual models outperform interleaved ones on all three metrics. Looking at the entire learning trajectory, interleaved models differ markedly from both monolingual and sequential models, showing more gradual and delayed improvements.
@@
This RQ concerned the CP for $\mathrm{L}_{1}$ attrition. Humans show few signs of $\mathrm{L}_{1}$ attrition after the CP, even if $\mathrm{L}_{1}$ exposure decreases or ceases. Our results in experiment 2 provide strong evidence against a similar phenomenon in typical LMs. Instead, we find that $\mathrm{L}_{1}$ performance worsens rapidly and to a large degree in the sequential condition after $\mathrm{L}_{1}$ exposure ceases. This is expected given the susceptibility of neural networks to catastrophic forgetting. The loss of $\mathrm{L}_{1}$ proficiency is prevented by continuing $\mathrm{L}_{1}$ exposure in the sequential-interleaved condition, but such continued exposure is not necessary in humans. [^15] This ignores self-talk as a potential source of $\mathrm{L}_{1}$ exposure; self-talk may continue even if external $\mathrm{L}_{1}$ exposure ceases.
@@
This RQ concerned the trade-off between $\mathrm{L}_{2}$ learning and preventing $\mathrm{L}_{1}$ attrition when explicitly reducing plasticity partway during learning. We find strong evidence for such a trade-off when comparing the sequential-EWC condition to the sequential condition. The value of $\lambda$ we selected preserved $\mathrm{L}_{1}$ performance substantially compared to the sequential models, at the cost of harming final performance in $\mathrm{L}_{2}$. The $\mathrm{L}_{2}$ learning curves also converge relatively quickly when plasticity is reduced. Our exploration of different values of $\lambda$ (Fig. 5) shows that preserving $\mathrm{L}_{1}$ to monolingual levels harms $\mathrm{L}_{2}$ acquisition by roughly 1.5 nats of $\mathrm{L}_{2}$ performance, but we do not directly compare grammatical performance of our models to that of human late $\mathrm{L}_{2}$ learners. Thus, our attempt at reverse engineering shows a broadly human-like learning pattern when using EWC, but we cannot say quantitatively and at a high level of granularity whether the result is human-like.
@@
This RQ concerned the impact of language similarity on CP effects. Our results showed that the language family of $\mathrm{L}_{1}$ and its script has an impact on $\mathrm{L}_{2}$ learning in the expected way for only a subset of evaluations. Based on earlier findings from Papadimitriou and Jurafsky (2020) and Oba et al. (2023), we had expected $\mathrm{L}_{2}$ performance would be greater when $\mathrm{L}_{1}$ is more closely related. However, only our results for PPL support this prior conclusion. Our results for BLiMP do not unless EWC is applied, suggesting that models are ordinarily plastic enough to learn the grammar of $\mathrm{L}_{2}$ regardless of relatedness, unless plasticity is specifically reduced. We note that Papadimitriou and Jurafsky (2020) reduce plasticity by freezing the model weights before transferring to $\mathrm{L}_{2}$, but we do not venture an explanation why Oba et al. (2023) seemingly find models to be less plastic than we do.
@@ 7.2 Theoretical Implications
Our results show that CP effects are not naturally arising in LMs in a typical training regime. At the same time, we are able to suggest a methodology to reverse engineer human-like learning patterns by artificially reducing plasticity later in training. As discussed in Section 3, results like these are relevant to certain specific claims in the critical period literature. Specifically, they refute the strong experiential claim, which states that all successful learning algorithms will show CP effects, and they are consistent with (but do not provide strong evidence for) the strong innate claim, which states that innate maturational stages are necessary to produce CP effects.
The strong experiential claim incorrectly predicts that our LMs will naturally show CP effects. This may come as a surprise, given this claim was based on previous studies on connectionist models finding evidence for the phenomenon of entrenchment (Munro, 1986; Ellis and Lambon Ralph, 2000). One explanation for this discrepancy in results is the difference in models’ capacity. These earlier works trained extremely small models by today’s standards, whereas our LMs may be over-parameterized and therefore have sufficient capacity to train to convergence without entrenchment.
On the other hand, the strong innate claim correctly predicts that our LMs will not show CP effects, unless we introduce an innate loss in plasticity. Our introduction of EWC part-way through training is akin to an innate loss in plasticity. However, our experiments are not strong evidence either that the strong innate claim is correct or that EWC is a plausible algorithmic mechanism underlying the loss in plasticity in humans. The possibilities remain that other statistical learners will show CP effects as a natural consequence of experience, and that humans could be one such learner.
Finally, many scholars of human development advocate for a complex explanation of CP effects involving both innate and experiential mechanisms (e.g.,  Newport, 1990; Thiessen et al., 2016; Singleton, 2005). We consider this nuanced view to be likely correct, but our results suggest that the role of statistical learning should not be assumed or overstated without evidence from humans or more cognitively plausible models.
@@ 7.3 Limitations and Future Work
From a Bayesian epistemological point of view, results from GPT-2 and RoBERTa should affect our priors about general learners and humans in some ways, but Transformer LMs are inherently limited as models of human learners. Assuming a sort of Copernican Principle for cognitive modeling, humans and LMs should both be unextraordinary relative to the theoretical class of language learners. So without other evidence or a priori reasoning, we should assume they share properties: i.e., the property that we identified in LMs that the ordinary mechanism of learning fails to lead to CP effects. However, there are many differences that could lead us a priori to hypothesize differences in how humans and LMs learn. More generalizable evidence could be obtained by considering more cognitively-inspired models and learning algorithms, and learning environments drawing on more developmentally plausible linguistic data and multimodal input (Warstadt and Bowman, 2022). For those who want to defend the position that CP effects fall out in humans—but not LMs—from ordinary learning, these or other differences must be identified as the cause. Our experiments seek a compromise given today’s resources by choosing architectures and procedures that maximize human-like learning outcomes, while still using a transcribed-speech training corpus and human-scale data. Future work should revisit this compromise as better resources for cognitive modeling are developed.
One caveat to the rejection of strong experiential claim is that the claim only applies to learners that adequately acquire $\mathrm{L}_{1}$. While this condition is not rigorously defined, one might argue our models do not qualify. For example, our GPT-2 models in the sequential and interleaved conditions achieve $75\pm 1\%$ accuracy on BLiMP. This is substantially better than chance (50%) and is comparable to the strongest baseline model from the BabyLM Challenge (Warstadt et al., 2023), but falls well below human agreement with BLiMP (89%). No LMs currently achieve fully human-like performance on BLiMP, and higher performance generally requires compromising on developmental plausibility, though more data-efficient architectures exist (Samuel et al., 2023; Warstadt et al., 2023). We do not expect our results to be qualitatively different if our experiments are run with more effective Transformer-based LMs, but this is something future work should confirm.
@@
We identify several additional avenues for future work: First, the regularization method we use, EWC (Kirkpatrick et al., 2017), can be viewed only as a computational-level model of an innate biological CP; future work could consider other regularizers such as Memory Aware Synapsis (Aljundi et al., 2018) that arguably model what happens in humans at the algorithmic level. Second, there is neurolinguistic evidence for some degree of modularity in how human bilinguals process different languages (Hernandez et al., 2005), but our models use completely shared parameters for languages. Future work can explore LM architectures that encourage or directly build in modularity, such as XLM (Conneau and Lample, 2019) or X-MOD (Pfeiffer et al., 2022). Third, our models learn in a text-only environment, but for humans $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ are both grounded in the same non-linguistic stimuli. Training multimodal models can lead to more realistic simulations, as well as enable testing of CP effects for $\mathrm{L}_{1}$ learning, which requires non-linguistic experience to precede $\mathrm{L}_{1}$ acquisition.
@@ 8 Conclusion
Acknowledgments
We thank the editors and reviewers whose feedback has led to substantial improvements in this work. TP and AW are supported by ETH Postdoctoral Fellowships.
@@ Appendix A Data Preprocessing
We remove extra spaces, non-breaking spaces and dataset-specific characters such as dialogue lines and music symbols (in OpenSubtitles) and paragraph delimiters (in Gutenberg). For the Java code extracted from The Stack, we remove all docstrings and comments.
@@
Interleaved datasets are created for each language pair ($\mathtt{en}$ plus a second language). The same blocks of texts sampled in the previous unifying step are simply interleaved while maintaining their ordering (e.g., $\mathrm{L}_{1}$ block 1, $\mathrm{L}_{2}$ block 1, $\mathrm{L}_{1}$ block 2, $\mathrm{L}_{2}$ block 2, etc.). In this way, the data from the two languages is presented in the same order to the model during the interleaved training as it is during the sequential training.
@@
We uniformize the training dataset sizes across languages. We quantify dataset sizes in terms of the number of tokens obtained using a BPE tokenizer; see section App. A. Beyond the fact that tokens are the true unit of input to the LMs, BPE is a compression algorithm, so while the amount of information per word might be highly language-specific, the amount of information per token is more comparable across languages.
@@ Appendix C Hyperparameters
Table 4: Configurations for GPT-2 models.
@@
Table 5: Configurations for RoBERTa models.
@@
Table 6: Fixed hyperparameters for GPT-2 and RoBERTa models.
@@ Appendix D Evaluation Tasks
Irregular Forms Aaron broke the unicycle. Aaron broken the unicycle.
@@
Ellipsis Jill hides one orange chair and Tammy hides more. Jill hides one chair and Tammy hides more orange.
@@
The BLiMP tasks that our models are evaluated on are examplified in Tab. 7. The subset of GLUE  [^17] https://gluebenchmark.com tasks that our models are evaluated on are the following.
@@
The Corpus of Linguistic Acceptability consists of around 10K English sentences gathered from published linguistics literature that have been annotated by experts for binary acceptability (grammaticality) judgements (Warstadt et al., 2019). The evaluation metric is Matthews Correlation Coefficient.
@@
The Binary Stanford Sentiment Treebank includes 215K unique phrases extracted from the parse trees of movie reviews sentences and that have been fully labeled as having either a positive or negative sentiment by three human judges (Socher et al., 2013). The evaluation metric is Accuracy.
@@
The Microsoft Research Paraphrase Corpus contains 5,800 pairs of sentences sourced from a large corpus of news data, each labeled with a binary judgment indicating whether the pair represents a paraphrase or not (Dolan and Brockett, 2005). The evaluation metric is F1 score.
@@
The Quora Question Pairs is a corpus of over 400K question pairs from the Quora website which are annotated with a binary value denoting whether the questions are paraphrases of each other. The evaluation metric is F1 score.
@@
The Multi-Genre Natural Language Inference is a crowd-sourced collection of 433K sentence pairs annotated with textual entailment information, covering a wide range of genres of both spoken and written text (Williams et al., 2018). The evaluation metric is Accuracy.
@@
The Multi-Genre Natural Language Inference Mismatched dataset is the mismatched version of the MNLI (matched) dataset, where the dev and test sets use out-of-domain data that does not closely resemble anything seen at training time (Williams et al., 2018). The evaluation metric is Accuracy.
@@
The Question-answering Natural Language Inference dataset is automatically derived from the Stanford Question Answering Dataset v1.1 (SQuAD) (Rajpurkar et al., 2016). It consists of question–paragraph pairs with one sentence in each paragraph, sourced from Wikipedia, containing the answer to the corresponding question, written by an annotator (Wang et al., 2019b). The evaluation metric is Accuracy.
@@
The Recognizing Textual Entailment dataset is compiled from a series of textual entailment challenges: RTE1 (Dagan et al., 2005), RTE2 (Bar-Haim et al., 2006), RTE3 (Giampiccolo et al., 2007), and RTE5 (Bentivogli et al., 2009). The task requires to recognize whether the meaning of a text fragment can be inferred from the other text. The evaluation metric is Accuracy.
The subset of SuperGLUE  [^18] https://super.gluebenchmark.com tasks that the models are evaluated on are the following.
@@
The Boolean Questions dataset is a reading comprehension dataset that consists of almost 16K naturally occurring yes–no questions generated in unprompted and unconstrained settings (Clark et al., 2019). The evaluation metric is Accuracy.
@@
The Multi-Sentence Reading Comprehension dataset contains around 10K questions that can be answered by combining information from a multi-sentence paragraph (Khashabi et al., 2018). The evaluation metric is F1 score.
@@
The Winograd Schema Challenge dataset is a corpus of sentence pairs that differ in only one or two words and contain an ambiguity that can be resolved using world knowledge and reasoning (Levesque et al., 2011). The evaluation metric is Accuracy.
Appendix E Derivation of EWC Regularization for Language Modeling
Let $\Sigma$ be an alphabet; furthermore, define $\overline{\Sigma}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\Sigma\cup\{\textsc{eos}\}$. We assume we have access to a collection of strings $\mathcal{D}_{\mathrm{L}_{1}}=\{\boldsymbol{x}^{(n)}\}_{n=1}^{N}\subset\Sigma^{*}$ in $\mathrm{L}_{1}$ and another $\mathcal{D}_{\mathrm{L}_{2}}=\{\boldsymbol{y}^{(m)}\}_{m=1}^{M}\subset\Sigma^{*}$ in $\mathrm{L}_{2}$. Additionally, let $\Theta\subset\mathbb{R}^{d}$ be a compact set of possible parameters. We take a Bayesian approach and construct a posterior density over possible parameter vectors $\boldsymbol{\theta}$. Let ${\pi}(\boldsymbol{\theta})$ be a prior density over $\Theta$, and consider the following likelihood
$p(\mathcal{D}_{\mathrm{L}_{1}},$ $\mathcal{D}_{\mathrm{L}_{2}}\mid\boldsymbol{\theta})=\underbrace{\prod_{n=1}^{N}p(\boldsymbol{x}^{(n)}\mid\boldsymbol{\theta})}_{\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}p(\mathcal{D}_{\mathrm{L}_{1}}\mid\boldsymbol{\theta})}\underbrace{\prod_{m=1}^{M}p(\boldsymbol{y}^{(m)}\mid\boldsymbol{\theta})}_{\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}p(\mathcal{D}_{\mathrm{L}_{2}}\mid\boldsymbol{\theta})}$ (2a)
$=\prod_{n=1}^{N}\prod_{t=1}^{T_{n}}p(\textsc{eos}\mid\boldsymbol{x}^{(n)},\boldsymbol{\theta})p(x_{t}^{(n)}\mid\boldsymbol{x}_{<t}^{(n)},\boldsymbol{\theta})\prod_{m=1}^{M}\prod_{t=1}^{T_{m}}p(\textsc{eos}\mid\boldsymbol{y}^{(m)},\boldsymbol{\theta})p(y_{t}^{(m)}\mid\boldsymbol{y}_{<t}^{(m)},\boldsymbol{\theta}),$ (2b)